OpenStudy (anonymous):

How many solutions does the system have? x = -4y + 4 2x + 8y = 8

OpenStudy (amorfide):

if you rearrange it into the same format as the other equation either ax+by=c or ay=bx+c so if you have x = -4y + 4 2x + 8y = 8 and rearrange to x + 4y =4 2x + 8y = 8 now take the same steps as you would for solving simultaneous equations multiply equation 1 by 2 to get 2x+8y=8 2x+8y=8 you can see they are the same line so this means they have infinitely many solutions